OpenStudy (brittyarellano):

Using the following formula for continuous compounding interest, A=Pe^rt, calculate the amount of interest a credit card with an annual interest rate of 29%would generate in a month if you had a starting amount of 11,000 I just would like to know how to set up this equation

OpenStudy (wolf1728):

In that equation A is final amount P is beginning amount e = 2.718281828 r is the interest rate t is the time (month would be 1/12 = 0.083333333) final amount = 11,000 * 2.718281828^(.29*0.083333333) The final amount will be the interest plus the principal.

OpenStudy (brittyarellano):

but its asking how much the annual interest rate would generate in a month

OpenStudy (brittyarellano):

how would i find that

OpenStudy (wolf1728):

r = .29 annual rate r = .29/12 = 0.0241666667 monthly